OpenStudy (anonymous):

Write the exponential equation in logarithmic form. x0 = 1, (x is different than 0)

OpenStudy (experimentx):

?? ... looks like you misunderstood question

OpenStudy (anonymous):

x^0 = 1, (x is different than 0)

OpenStudy (experimentx):

y = b^x is an exponential equation, where b is a constant the logarithmic form would be logb y = x where log is logb base b. for your question, logx1 = 0